### v1

# Multi-directional investigations on quiet time suprathermal ions measured by ASPEX-STEPS on-board Aditya L1

Abstract
The origin, acceleration and anisotropy of suprathermal ions in the interplanetary medium during quiet periods have remained poorly understood issues in solar wind physics. To address these aspects, we derive the spectral indices for the “quiet” time suprathermal ions based on the measurements by the four directionally separated sensors that are part of the Supra-Thermal and Energetic Particle Spectrometer (STEPS) of Aditya Solar Wind Particle EXperiment (ASPEX) on-board Aditya L1 spacecraft. Three out of four STEPS sensors - Parker Spiral (PS), Inter-Mediate (IM), Earth Pointing (EP) - are in one plane (nearly aligned with the ecliptic plane) while the fourth sensor - North Pointing (NP) – is in a mutually orthogonal plane. The energy ranges covered by the PS, IM, EP and NP sensors are 0.36-1.32 MeV, 0.14-1.22 MeV, 0.39-1.33 MeV and 0.12-1.23 MeV respectively. The quiet intervals are identified during January - November, 2024 and the derived spectral indices (differential directional flux versus energy) are found to be in the range of 2.0 for all directions in the time scale of a few days revealing isotropic nature of their distribution. Further analysis of elemental abundance ratios (${}^{3}\mathrm{He}/{}^{4}\mathrm{He}$, $\mathrm{Fe}/\mathrm{O}$, and $\mathrm{C}/\mathrm{O}$) during the same quiet intervals obtained from the Ultra-Low Energy Isotope Spectrometer (ULEIS) on –board the Advanced Composition Explorer (ACE) spacecraft suggests possible contributions from the leftover ions from the previous impulsive (Solar flares) and gradual events (CMEs) in the “quiet” time suprathermal ion pool.
[…]
show]aakashgupta.du@gmail.com
1 Introduction
An important feature of the solar wind is the suprathermal ion tail which is manifested in the form of ions in the energy range of $\sim 10~{}\mathrm{keV~{}n^{-1}}$ to $\sim 1~{}\mathrm{MeV~{}n^{-1}}$. These suprathermal ions are ubiquitous in the heliosphere and are continuously observed to arrive at the spacecraft location from multi-directions in the interplanetary (IP) medium. The suprathermal ions act as a seed population for solar energetic particles (SEPs) accelerated by IP shocks associated with transient events in the heliosphere such as coronal mass ejections (CMEs; J. Gosling et al. 1981; G. Mason et al. 1999; M. I. Desai et al. 2003, 2004) and corotating interaction regions (CIRs; L. Fisk & M. Lee 1980; K. Chotoo et al. 2000; H. Kucharek et al. 2003; G. M. Mason et al. 2008; R. C. Allen et al. 2019). In astrophysical plasmas, particle acceleration is primarily governed by two well-established mechanisms: first-order Fermi acceleration, also known as diffusive shock acceleration (G. Krymskii, 1977; A. Bell, 1978), and second-order Fermi acceleration, originally proposed by E. Fermi (1949). Both mechanisms require the existence of a pre-accelerated seed ion population whose energies are above the injection threshold and which can be further energized within the acceleration framework. Therefore, the importance of suprathermal ion pool is undeniable to understand the generation of SEPs.
The IP medium hosts a diverse reservoir of suprathermal ions, consisting of protons as well as heavier ions (4He to Fe and beyond) (M. Desai & J. Giacalone, 2016a). Studies of the composition of these suprathermal ion populations indicate that their origins can be traced to multiple sources. These sources include solar wind (e.g., M. I. Desai et al. 2003), remnants from the past transient events such as CMEs and solar flares (G. Mason et al., 1999; D. V. Reames, 2013), and interstellar pick-up ions (R. C. Allen et al., 2019). Notably, beyond 1 AU, interstellar pick-up ions have been found to contribute into the suprathermal ion pool (L. Fisk, 1976). Their abundance and energy distribution play a crucial role in shaping the overall dynamics of suprathermal ion population and contribute to the seed population for the subsequent shock acceleration in the heliosphere (A. Tylka et al., 2001). Suprathermal ions typically exhibit power-law spectra when the differential directional flux (phase space density) is plotted as a function of energy (speed). In general, a power law spectrum and its spectral index (slope of the power spectrum) correspond to a natural stochastic process. In case of suprathermal particles also, theoretical models such as pump acceleration mechanism suggests that this power-law spectrum follows a characteristic spectral index of approximately $-1.5\ (-5)$ in the differential directional flux vs. energy (phase space density vs. speed) representation (L. Fisk & G. Gloeckler, 2006, 2008, 2012, 2014) during “quiet” times. However, several studies (e.g. M. Dayeh et al. 2009, 2017; B. Dalal et al. 2022) show that the spectral index of suprathermal ions in the 100 keV to 1 MeV energy range during “quiet” times is not universally fixed at $-1.5$, but varies over a wide range (1.22.91.22.91.2–2.91.2 – 2.9 ). Moreover, these works also reveal that, on many occasions, different elements have significantly different spectral indices during a given interval suggesting $m/q$ dependence as far as acceleration mechanisms are concerned (e.g. B. Dalal et al. 2022). Therefore, the understanding on the source and energization process(es) of suprathermal ions in the IP medium is far from complete.
While dealing with acceleration of suprathermal ions in the interplanetary (IP) medium, Parker transport equation (PTE) is solved on many occasions with the assumption that suprathermal ions are isotropic in the IP medium (e.g. L. Fisk & G. Gloeckler 2008). Since suprathermal ion fluxes obtained from the Supra-Thermal and Energetic Particle Spectrometer (STEPS) of Aditya Solar Wind Particle EXperiment (ASPEX) on-board India’s Aditya L1 are directionally resolved, experimental verification of the isotropic assumption is possible based on evaluation of spectral indices. Further, unlike previous works (e.g. M. Dayeh et al. 2009, 2017; B. Dalal et al. 2022) where datasets for longer intervals are used, we use shorter intervals of data from ASPEX-STEPS for this study to minimize influence of multiple transient processes in the suprathermal ion fluxes. After selecting shorter intervals (a few days) of quiet period fluxes in 2024, we check the variations in spectral indices with directions.
It is to be noted here that although directionally resolved fluxes are obtained from ASPEX-STEPS, these fluxes are not species separated and are primarily dominated by protons and alpha particles. Therefore, in order to understand the source process(es) through which these quiet period suprathermal ions are generated, elemental abundance data for the same quiet intervals are analyzed. We use the data from Ultra Low Energy Isotope Spectrometer (ULEIS) on –board the Advanced Composition Explorer (ACE) spacecraft for this purpose.
Therefore, we intend to address two problems in this work. One, whether the measured suprathermal ion fluxes are directionally isotropic at L1 and two, what are the possible sources of these ions. The manuscript is organized as follows. The datasets are discussed in section 2. In Section 3, the methodology used to select the quiet intervals is discussed. The salient results are presented in Section 4. In Section 5, we discuss the results and in Section 6, the salient inferences from this work are summarized.
2 Dataset
Aditya L1 (S. Seetha & S. Megala, 2017; D. Tripathi et al., 2022) is the first observatory class mission of India to study the sun and solar wind. Aditya L1 was launched on 2nd September 2023 and placed in the halo orbit around the first Lagrange point (L1) of the Sun-Earth system on 6th January 2024. The Aditya Solar wind Particle EXperiment (ASPEX) (P. Janardhan et al., 2017; S. Goyal et al., 2018; P. Kumar et al., 2025) is one of the three in-situ experiments that measures the solar wind and energetic ions. ASPEX consists of two subsystems and the Supra-Thermal and Energetic Particle Spectrometer (STEPS) (S. K. Goyal et al., 2025) subsystem of ASPEX measures the suprathermal and energetic ions. There are six sensors in STEPS – Sun-Radial (SR), Inter-Mediate (IM), Parker Spiral (PS), North Pointing (NP), South Pointing (SP) and Earth Pointing (EP). Among the six sensors, the SR and SP sensors are experiencing light saturation issues and data from these sensors are not used in this work. As described in S. K. Goyal et al. (2025), the NP and IM units are a single-window sensor equipped with a dead layer of $0.2\,\mu\mathrm{m}$ thickness. In contrast, the PS and EP units consist of a stack of custom-built dual-window Si-PIN detector and a combined scintillator plus silicon photomultiplier (SiPM) detector assembly. The dual-window Si-PIN detector features two active regions with different dead layer thicknesses: $0.1\,\mu\mathrm{m}$ for the circular inner detector (PS-Inn and EP-Inn, with a diameter of $7\,\mathrm{mm}$) and $0.8\,\mu\mathrm{m}$ for the annular outer detector (PS-Out and EP-Out, spanning diameters from $7\,\mathrm{mm}$ to $18\,\mathrm{mm}$). Due to the extremely thin dead layer in the PS-Inn and EP-Inn sensor, it exhibits distinct spectral signatures corresponding to multiple ion species. As the deconvolution of these spectral features in the PS-Inn and EP-Inn sensor is still ongoing, this study utilizes species-integrated energetic ion flux data from the PS-Out and EP-Out sensors. Consequently, references to the PS and EP unit in this work pertain exclusively to the PS-Out and EP-Out sensors. The fluxes measured by the STEPS sensors are converted from the spacecraft frame to the GSE coordinate system. The three (PS, IM and EP) out of four remaining sensors measure primarily in the X-Y plane of the GSE coordinate system whereas the NP sensor is directed toward the +ve Z direction. STEPS measures the flux of suprathermal/energetic ions in the energy range spanning from 20 KeV/n to 6 MeV/n. We use the linear (in a log-log curve of flux vs. energy) part of the spectra to derive the spectral slopes from the fluxes. This leads to selection of ion fluxes with lower cut-off of 0.3 MeV for PS and EP sensors and 0.12 MeV for IM and NP sensors. Further, fluxes above 1.3 MeV for PS and EP or above 1.2 MeV for IM and NP are also not used for two reasons. One, fluxes in the higher energy channels are negligible during quiet periods and two, we avoid solar energetic particles ($>1.5\,\mathrm{MeV}$) in this work. For elemental abundance of suprathermal ions, data of ${}^{3}\mathrm{He}$, ${}^{4}\mathrm{He}$, C, Fe and O from the Ultra-Low Energy Isotope Spectrometer (ULEIS; E. C. Stone et al. 1998) on board the Advanced Composition Explorer (ACE) spacecraft are used. Since the objective of this work is to investigate quiet periods, the criterion for selection of quiet periods is described in the subsequent section.
3 Methodology
This section is divided into two subsections. The first subsection deals with the methodology based on which the quiet intervals are chosen and the second subsection contains the method on how spectral slopes are calculated.
3.1 Selection of quiet intervals
By quiet time, we refer to the conditions in the IP medium when there is no enhancement in the suprathermal ion fluxes beyond a threshold level due to the effects of transient events like SIR/CIRs, ICMEs etc. Quiet times are selected following a heuristic statistical methodology suggested earlier by M. Dayeh et al. (2017). In this method, as a first step, hourly averaged directional differential flux is calculated based on the measurements by AL1-ASPEX-STEPS-PS, IM, EP, and NP sensors in the energy ranges of 0.36 – 1.32 MeV, 0.13 – 1.22 MeV, 0.38 – 1.32 MeV, and 0.12 – 1.23 MeV, respectively. Subsequently, the fluxes from each direction are sorted in an ascending order. These hourly-averaged and sorted fluxes are re-binned with bin size of 24 hours and the mean value as well as variance are calculated. The mean versus variance values of sorted fluxes for each sensor of AL1-ASPEX-STEPS are plotted in Figure 1.
Figure 1: Mean versus Variance plots of the sorted (in ascending order) directional differential flux for Parker Spiral (PS), Earth Pointing (EP), North Pointing (NP), and Intermediate (IM) sensors of AL1-ASPEX-STEPS during the observational period from January to November 2024. The vertical red dashed line in each panel represents the threshold value of the flux below which the fluxes are considered as the quiet time flux and beyond which the fluxes are considered to be associated with transient events. The threshold values are also mentioned. The corresponding energy range is written at the top of each panel.
The vertical dashed red line in each subplot corresponds to the threshold value of the mean flux beyond which the variance in flux increases monotonically. It is apparent that the mean flux values beyond the threshold value represent enhanced fluxes during transient events (CMEs, CIRs/SIRs). Consequently, the mean flux values below the threshold value represent quiet time flux of suprathermal ions in the IP medium. The quiet intervals identified for each of the detectors for the month of March 2024 is shown in Figure 2.
Figure 2: Panels a, e, c and g present the temporal variations of differential directional fluxes of suprathermal ions measured by PS, IM, EP, and NP detectors of AL1-ASPEX-STEPS, respectively. Panels b, f, d, and h present the corresponding quiet time directional differential flux of suprathermal ions in PS, IM, EP, and NP detectors, respectively. The energy channels for the panels a, b, c, d and e, f, g, h are mentioned at the left side and right side of these panels respectively.
In Figure 2, the original variation (panels a, e, c, g) in fluxes is juxtaposed with the quiet time variations (panels b, f, d, h) to show the typical quiet intervals. This methodology is applied for the period from January 2024 to November 2024 except for the month of April 2024 when the satellite underwent many rotations and thruster firings. Spectral indices of suprathermal ions have been calculated for all the quiet intervals obtained using the above-mentioned methodology and this is discussed in the next subsection.
3.2 Estimation of spectral index
We present below (Figure 3) typical examples of ion spectra obtained from PS, IM, EP and NP sensors.
Figure 3: Plots of quiet time differential directional flux versus energy with spectral indices value mentioned in each subplot for the PS (a), IM (b), EP (c) and NP (d) directions of AL1-ASPEX-STEPS. Spectral indices are calculated by fitting only the linear part in this log-log plots. The range of energy for the fitted linear part is mentioned in each subplot for PS, IM, EP and NP in red, blue, green and magenta colors respectively.
In figure 3, we show the typical Ion spectra for PS, IM, EP and NP sensors during 07 March, 2024 to 10 March 2024. As can be seen, the nature of the spectra is different at the lower energies. This is due to low-level discriminator (LLD) cut-off at the instrument level. This is due to low-level discriminator (LLD) threshold, which is set electronically to reduce the instrument background noise. The counts below this LLD threshold are considered to be zero and above this value, there is non-linear behaviour up to a certain energy as seen in Figure 3. We fit $J=AE^{-m}$ function to the linear portions of the spectra to calculate the spectral indices. The linear portion of the spectra (in log-log curve) for PS, IM, EP and NP sensors are found to be within the energy ranges of 0.36-1.32 MeV, 0.14-1.22 MeV, 0.39-1.33 MeV and 0.12-1.23 MeV respectively. The energy bins are calculated within these linear fit regions. Thereafter, spectral indices along with the standard errors (fitting errors) are calculated using these energy bins. In this analysis, it is to be noted that the counting uncertainties are also considered while estimating the spectral indices. The counting uncertainties in the flux values are shown as vertical error bars. Since the vertical error bars are very small, these are not clearly visible in log-log scale.
4 Results
In this section, the spectral indices are calculated for all the four ASPEX-STEPS sensors (PS, IM, EP, and NP) under three different scenarios. As the quiet intervals are not always identical for all the four sensors, in section 4.1, we find out common quiet intervals and derive the spectral indices for the four sensors. In section 4.2, we focus on individual sensors and consider all the quiet intervals for that sensor. Further, we also consider a period when the Aditya-L1 spacecraft was rotated for some time and therefore, the four STEPS sensors also rotated from their nominal positions during this interval. This scenario (section 4.3) is brought in to verify whether the spectral indices derived in the first two scenarios remain consistent in the third scenario. The spectral indices derived for these three scenarios are presented in the ensuing subsections.
4.1 Common quiet intervals for PS, IM, EP and NP sensors
We present in Figure 4 the spectral fits and the corresponding indices calculated for a few common quiet intervals recorded by the PS, IM, EP and NP sensors. The duration of these intervals ranges from a day to a few days as can be seen from the top of each of the subplots. Each row in Figure 4 represents a particular sensor but for 5 different intervals. The values of spectral indices (standard fitting errors are also mentioned) for common quiet intervals are for all the sensors are found to be nearly similar, i.e. $\sim 2.0$, suggesting nearly identical spectral indices in all directions.
Figure 4: Plots of quiet time differential directional flux versus energy with spectral indices value mentioned in each subplot for the PS (a-e), IM (f-j), EP (k-o) and NP (p-t) sensors of AL1-ASPEX-STEPS.
4.2 Quiet intervals for individual sensors (PS, IM, EP and NP)
In this section, we present (Figures 5-8) the spectral fits as well as indices calculated for the quiet intervals selected for each individual sensor. Similar to Figure 4, here also the quiet intervals are mentioned at the top of each sub-plots. The spectral indices calculated for quiet intervals of each sensor also show similar values of $\sim 2.0$. Therefore, regardless of the intervals and sensors, the spectral indices are found to be close to 2.
Figure 5: Plots of quiet time differential directional flux versus energy with spectral indices value mentioned in each subplot for the Parker Spiral (PS) sensor of AL1-ASPEX-STEPS.
Figure 6: Plots of quiet time differential directional flux versus energy with spectral indices value mentioned in each subplot for Intermediate (IM) sensor of AL1-ASPEX-STEPS.
Figure 7: Plots of quiet time differential directional flux versus energy with spectral indices value mentioned in each subplot for Earth pointing (EP) sensor of AL1-ASPEX-STEPS.
Figure 8: Plots of quiet time differential directional flux versus energy with spectral indices value mentioned in each subplot for North pointing (NP) sensor of AL1-ASPEX-STEPS.
4.3 Quiet intervals when AL1 was rotated (during November - 2024)
During the quiet interval of 25 November 2024, Aditya L1 (AL1) spacecraft underwent a rotation for a few hours (06:30 UTC to 14:10 UTC). In this rotation period, PS, IM, EP and NP detectors also rotated from their nominal positions. Through Figure 9, we illustrate the orientations of the PS, IM, NP, and EP sensors in their nominal and rotated configurations, represented in the Geocentric Solar Ecliptic (GSE) coordinate system. Panels (a–h) show the nominal configuration, while panels (i–p) display the rotated configuration. In all panels, the X-axis represents the direction cosine corresponding to the $X_{GSE}$ component for each of the four sensors. Panels (a–d) and (i–l) present the direction cosines of the $Y_{GSE}$ component along the Y-axis, whereas panels (e–h) and (m–p) show the direction cosines of the $Z_{GSE}$ component along the Y-axis. Each panel includes a blue and a yellow dot denoting the Earth and the Sun, respectively. The red and green arrows indicate the orientations of the PS, IM, NP, and EP sensors projected in the X–Y and X–Z planes of the GSE coordinate system, respectively. The length of each arrow is proportional to the corresponding direction cosine component, thus accurately reflecting the detector orientation.
Figure 9: Direction cosines of PS, IM, NP and EP sensors in nominal configuration (panel a-h) and in rotated configuration (panel i-p) of AL1-ASPEX-STEPS. In subplots (a-p) blue dot and yellow dot represents the Earth and the Sun respectively. Red and green arrows represent the directions of PS, IM, NP and EP sensors in XY and XZ plane of GSE coordinate system respectively. The length of an arrow is decided by the corresponding values of direction cosines.
We estimated the spectral indices for the PS, IM and NP sensors during this quiet interval when the spacecraft was rotated (shown in Figure 10). For the EP sensor, the quiet time criteria did not get satisfied and hence spectral indices are not calculated for the EP sensor during this period. Interestingly, spectral indices calculated during this interval also show the similar values $\sim 2.0$ regardless of the sensors and directions.
Figure 10: Plots of quiet time differential directional flux versus energy with spectral indices value mentioned in each subplot for PS, IM and NP sensors of AL1-ASPEX-STEPS during rotated period.
In summary, the Figures 4-10 reveal that spectral indices are close to 2 regardless of the direction and sensor during quiet intervals.
5 Discussions
In this study, we have calculated spectral indices for quiet time suprathermal ions in IP medium using measurements from the four sensors of STEPS instrument of ASPEX onboard Aditya L1 spacecraft for different quiet intervals as well as during nominal and rotated configurations of the spacecraft. The value of spectral indices obtained from PS, IM, EP and NP sensors are found to be $\sim 2.0$ that suggests isotropic distribution of quiet time suprathermal ions. A few earlier studies (e.g. M. Dayeh et al. 2009, 2017; B. Dalal et al. 2022) have shown that the spectral indices vary over a wide range even for quiet periods. These studies were based on the longer intervals of data obtained from instruments onboard spinning spacecraft such as Wind and Advanced Composition Explorer (ACE). Our analysis is done for shorter intervals of measurements in a directionally resolved manner. In order to assess whether this apparent isotropy reflects a genuine physical property of the ST ions distribution or is influenced by the relative motion between the spacecraft and the solar wind frame, a correction was applied using the Compton-Getting effect (F. Ipavich, 1974). The Compton-Getting effect accounts for the alteration in the observed energy spectra of particles due to the Doppler-like shift arising from the motion of the observer with respect to the solar wind rest frame. This correction enables a more accurate inference of the intrinsic particle distribution by eliminating systematic effects related to the observer’s velocity. After considering the Compton-Getting effect, we obtain the corrected or effective spectral index, which better represents the true properties of the particles.
$m_{\text{eff}}=m^{\prime}+\frac{(m^{\prime}+1)\left(\frac{w}{v}\right)\left[\cos\theta-\left(\frac{w}{v}\right)\right]}{1-2\left(\frac{w}{v}\right)\cos\theta+\left(\frac{w}{v}\right)^{2}}$ (1)
Where,
$m_{\text{eff}}$ = Spectral index in the observer’s frame (Spacecraft frame)
$m^{\prime}$ = Spectral index in the solar wind frame
$w$ = Relative speed between solar wind and spacecraft
$v$ = Speed of the particles/ions
$\theta$ = Angle between the solar wind velocity and the direction in which the observer is looking.
In this analysis, since the speed of suprathermal ions $\left(v\right)$ observed is much larger than the relative speed $\left(w\right)$ between solar wind and spacecraft, the term $\left(\frac{w}{v}\right)^{2}$ in the equation (1) becomes negligible. That’s why we applied the Compton- Getting effect up to the first order correction of $\left(\frac{w}{v}\right)$ in calculating the spectral indices. Equation (1) up to the first order of $\left(\frac{w}{v}\right)$ is given below.
$m_{\text{eff}}=m^{\prime}+\frac{(m^{\prime}+1)\left(\frac{w}{v}\right)\cos\theta}{1-2\left(\frac{w}{v}\right)\cos\theta}$ (2)
Spectral indices obtained by applying the correction due to the Compton- Getting effect do not exhibit the significant changes up to the first order and remains almost the same $\sim 2.0$. The reason behind the negligible changes in the spectral indices after applying Compton- Getting effect is that the energies (speeds) of the suprathermal ions are very high in comparison to relative speed $\left(w\right)$ between solar wind and spacecraft which makes the term $\left(\frac{w}{v}\right)$ negligible in equation 2. In our analysis, the angle $\theta$ between the solar wind velocity and the PS, IM, EP and NP direction is $52^{\circ}$, $30^{\circ}$, $156^{\circ}$, and $90^{\circ}$ respectively (for more details, refer S. K. Goyal et al. 2025). There will be no Compton-Getting correction in the North Pointing (NP) direction as $\theta=90^{\circ}$ which makes $\cos\theta$ term zero in equation 2. For all other angles, the Compton-Getting corrections are also minimal. Thus, negligible changes are seen in the values of spectral indices for PS, IM and EP directions after applying the Compton- Getting corrections. This shows that the isotropy is a genuine physical property of the suprathermal ions distribution and this is not influenced by the relative motion between the spacecraft and the solar wind frame. This investigation validates the assumptions of isotropic distribution of quiet time suprathermal ions made by L. Fisk & G. Gloeckler (2008) in solving Parker transport equation (PTE) for acceleration of suprathermal ions in the solar wind.
Another important insight that comes out of this study is that the spectral indices remain nearly identical $\left(\sim 2\right)$ for quiet intervals if we consider a time scale of a few days. This is in contrast to previous studies that reported significant variability in spectral indices when large intervals were considered. For instance, M. Dayeh et al. (2009, 2017) calculated spectral indices using a one-year average for eighteen years (1998-2015) and observed a wide range of values of spectral indices for different elements. Similarly, B. Dalal et al. (2022) analyzed spectral indices using a two-year average across different phases of the solar cycle—such as solar maximum, declining phase, solar minimum, and rising phase—and also reported substantial variations in the spectral index values for different elements. This significant variability in spectral index values indicates that even during the quiet period, different processes dominate in different phases of solar cycles. M. Dayeh et al. (2017) suggests that while impulsive and shock acceleration events dominate during enhanced solar activity periods, stream/corotating interaction regions (SIR/CIR) and heated solar wind processes dominate during low solar activity periods. These authors also suggest that the large variability of spectral indices above $100keV$ indicate towards the dominant role of these large, transient processes. Earlier studies by Y. Chen et al. (2015) and R. Hajra & J. V. Sunny (2022) have shown that SIR/CIRs are more prevalent during the declining and minimum phases of the solar cycle. B. Dalal et al. (2023) analyzed the spectral indices of suprathermal ions associated with multiple SIR/CIRs and found typical values of spectral indices in the range of 3-4, indicating significantly softer spectra. In contrast, our study, conducted during a solar maximum period, reports spectral indices closer to 2 in all the four directions. This suggests that during solar minimum, the quiet time suprathermal ion pool is primarily influenced by ions accelerated by CIRs/SIRs, resulting in softer spectra while during solar maximum, the suprathermal ion population during quiet periods is possibly dominated by large, impulsive and transient events like impulsive and gradual solar energetic particle events (SEPs), leading to harder spectral indices. In this context, the time scale of a few days is also important. We believe that the influence of the large transient processes is minimized by this choice leading to the nearly identical spectral indices recorded by all the four sensors of AL1-ASPEX-STEPS. As 2024 is a high solar activity year and AL1-ASPEX-STEPS will continue to make measurements, it will be interesting to see whether this nearly identical spectral index continue to hold during low solar activity period as well and if the magnitude of spectral index changes. Further, since ASPEX-STEPS does not discriminate between species and primarily measures protons and alpha particles (being the most abundant), the $m/q$ dependences of the acceleration processes are also averaged out. This suggests that it is difficult to comment on the individual processes that fix the spectral indices close to 2 at shorter time scales based on the present set of measurements. Therefore, in order to investigate the dominant processes influencing the isotropic distribution of quiet time suprathermal ions for the same periods as those of ASPEX-STEPS measurements, elemental abundance ratios were estimated using data from the Ultra-Low Energy Isotope Spectrometer (ULEIS) onboard the Advanced Composition Explorer (ACE) spacecraft. This is done following the same methodology suggested in the work of M. Dayeh et al. (2009, 2017).
Positioned at the L1 point, ACE is well-suited for this analysis as it provides species-resolved measurements of suprathermal ions. The elemental abundance ratios of ${}^{3}\mathrm{He}/{}^{4}\mathrm{He}$, $\mathrm{Fe}/\mathrm{O}$, and $\mathrm{C}/\mathrm{O}$ for the period from January to November 2024 are shown in Figure 11. Shaded regions in Figure 11 represent the same quiet periods for which spectral indices have been calculated using ASPEX-STEPS data. In Figure 11, black dots present the averaged elemental abundance ratios of ${}^{3}\mathrm{He}/{}^{4}\mathrm{He}$, $\mathrm{Fe}/\mathrm{O}$, and $\mathrm{C}/\mathrm{O}$ during quiet periods, with reference values associated with different particle populations—such as impulsive solar energetic particles (ISEPs) (G. Mason, 2007; D. V. Reames, 2021), gradual solar energetic particles (GSEPs) (G. Mason, 2007; D. V. Reames, 2021), corotating interaction regions (CIRs) (G. M. Mason et al., 2008), and the solar wind (J. Geiss & G. Gloeckler, 1998)—indicated by horizontal lines.
Figure 11: Plots of ${}^{3}\mathrm{He}/{}^{4}\mathrm{He}$, $\mathrm{Fe}/\mathrm{O}$, and $\mathrm{C}/\mathrm{O}$ ratios as measured by ULEIS/ACE in the energy range spanning from $0.1\,\mathrm{MeV/n}$ to $1.2\,\mathrm{MeV/n}$. The pink shaded regions represent the same quiet periods that are analyzed using the ASPEX-STEPS data. The horizontal lines represent averaged abundances measured in different heliosphere particle populations- CIRs abundance (CIRs; G. M. Mason et al. 2008); ${}^{3}\mathrm{He}$-rich or impulsive SEP events (impulsive SEPs; G. Mason 2007; D. V. Reames 2021); gradual SEP events (GSEPs) (gradual SEPs; G. Mason 2007; D. V. Reames 2021); Solar Wind (SW) value (J. Geiss & G. Gloeckler, 1998). Black dots in each subplot represent the average value of abundance ratios during quiet periods with one sigma error bar.
[…]
Table 1: Quiet time elemental abundance ratio of ${}^{3}\mathrm{He}/{}^{4}\mathrm{He}$, Fe/O, and C/O from ULEIS/ACE from 0.1 MeV/n to 1.2 MeV/n with standard deviation values.
Start date End date Avg 3He/4He Std 3He/4He Avg Fe/O Std Fe/O Avg C/O Std C/O
17-01-2024 04:00 21-01-2024 17:00 0.069 0.408 0.612 1.073 0.279 0.493
28-01-2024 19:00 28-01-2024 23:00 0.028 0.040 1.813 2.957 0.647 0.416
03-02-2024 09:00 03-02-2024 14:00 0.003 0.007 0.219 0.091 0.488 0.345
04-02-2024 12:00 06-02-2024 16:00 0.031 0.112 0.913 0.772 0.223 0.287
25-02-2024 18:00 29-02-2024 07:00 0.017 0.056 0.387 0.846 0.828 1.401
04-03-2024 00:00 15-03-2024 08:00 0.219 0.698 0.892 1.074 0.327 0.456
19-03-2024 01:00 23-03-2024 03:00 0.008 0.051 0.647 1.622 0.471 0.708
05-05-2024 22:00 08-05-2024 19:00 0.095 0.342 0.757 1.623 0.418 1.092
26-05-2024 15:00 30-05-2024 05:00 0.050 0.098 1.516 1.703 0.451 0.634
07-06-2024 01:00 08-06-2024 01:00 0.006 0.020 0.232 0.329 0.365 0.490
17-06-2024 21:00 24-06-2024 18:00 0.029 0.065 0.847 0.998 0.528 0.728
29-06-2024 16:00 30-06-2024 23:00 0.046 0.050 0.695 0.663 0.457 0.225
05-07-2024 11:00 17-07-2024 01:00 0.131 0.425 1.039 1.474 0.517 0.806
18-07-2024 12:00 21-07-2024 22:00 0.008 0.026 0.304 0.442 0.586 0.527
13-08-2024 00:00 15-08-2024 04:00 0.028 0.026 2.967 1.864 0.343 0.218
18-08-2024 04:00 29-08-2024 19:00 0.043 0.097 1.233 1.778 0.419 0.749
01-09-2024 02:00 01-09-2024 13:00 0.061 0.047 0.650 0.354 0.570 0.475
23-09-2024 15:00 25-09-2024 00:00 0.006 0.011 0.408 0.225 0.407 0.260
02-10-2024 03:00 04-10-2024 20:00 0.152 0.256 0.752 0.735 0.512 0.559
16-10-2024 15:00 24-10-2024 22:00 0.184 0.537 0.615 0.760 0.446 0.589
04-11-2024 00:00 05-11-2024 11:00 0.057 0.073 0.706 0.266 0.440 0.152
10-11-2024 09:00 13-11-2024 05:00 0.025 0.046 1.082 0.509 0.367 0.331
14-11-2024 05:00 21-11-2024 17:00 0.080 0.091 1.407 1.569 0.448 0.531
23-11-2024 16:00 30-11-2024 23:00 0.251 0.649 1.356 1.648 0.410 0.571
Insights drawn from Table 1 reveal notable signatures of impulsive solar energetic particle (ISEP) contributions during several intervals that are otherwise considered quiet. Specifically, substantial enhancements in the ${}^{3}\mathrm{He}/{}^{4}\mathrm{He}$ ratio commonly associated with ISEPs are observed during periods such as 04–15 March, 05–17 July, 02–04 October, 16–24 October, 14–21 November, and 23–30 November 2024. These elevated ${}^{3}\mathrm{He}/{}^{4}\mathrm{He}$ values, reaching or exceeding 0.1, suggest that approximately 25% of the selected quiet intervals are influenced by residual impulsive SEPs from previous events. In addition, the $\mathrm{Fe}/\mathrm{O}$ ratio exhibits considerable variability throughout the year. Notably, elevated $\mathrm{Fe}/\mathrm{O}$ values greater than 2, as seen on 28 January and during 13–15 August 2024, further support the presence of ISEP-like signatures. In contrast, approximately 45% of the quiet periods display $\mathrm{Fe}/\mathrm{O}$ ratios near unity (Fe/O $\sim 1$), indicating a substantial contribution from impulsive solar energetic particles (ISEPs) events, which are characterized by enhanced heavy ion abundances. The $\mathrm{C}/\mathrm{O}$ ratio adds further insight into the compositional makeup of the suprathermal ion population. Approximately 20% of the quiet intervals show $\mathrm{C}/\mathrm{O}$ values near 0.32, indicative of impulsive SEP contributions, while around 50% of the intervals display values near 0.42, which are characteristic of gradual SEP populations. Together, these findings underscore the complexity of particle composition even during nominally quiet times and highlight the persistent influence of impulsive events in shaping suprathermal ion populations.
6 Summary
In this study we focus on short-term quiet periods of suprathermal ions during solar maximum and the analysis leads to the following key findings:
  1. 1.
    The distribution of suprathermal ions during quiet periods is isotropic, as indicated by the similar spectral index values ($\sim 2$) observed in all four directions—Parker spiral (PS), intermediate (IM), Earth pointing (EP), and north pointing (NP).
  2. 2.
    This isotropy supports the assumption made by L. Fisk & G. Gloeckler (2008), who considered an isotropic suprathermal ion distribution while solving the Parker transport equation to explain the acceleration of suprathermal ions.
  3. 3.
    The suprathermal ion pool during quiet periods are predominantly composed of residual particles from previous impulsive and gradual solar energetic particles (SEPs) events in solar maximum conditions.
  4. 4.
    The ${}^{3}\mathrm{He}/{}^{4}\mathrm{He}$ abundance ratio indicates that approximately 25% of the quiet periods exhibit signatures of impulsive SEPs (ISEPs) contributions.
  5. 5.
    The $\mathrm{Fe}/\mathrm{O}$ ratio suggests that about 45% of the suprathermal ion population during quiet times can be attributed to ISEPs.
  6. 6.
    The $\mathrm{C}/\mathrm{O}$ ratio reveals that the suprathermal population comprises approximately 20% contribution from ISEPs and about 50% from gradual SEPs (GSEPs) events.
Authors also would like to thank various centres of ISRO for providing technical support and facilities for the test and calibration of ASPEX payload. Thanks are also due to the project, mission, and review teams of ISRO for their support. Authors are also grateful to the principal investigators and members of the ACE-ULEIS team for generating and managing the datasets used in this work. This work is supported by the Department of Space, Government of India. We greatly appreciate the discussion and suggestions given by Mihir Desai from Southwest Research institute (SWRI).
DATA AVAILABILITY STATEMENTS
The data used in this study are publicly available. The ASPEX-STEPS data from the Aditya-L1 mission are accessible from two sources: data from January 7 to May 10, 2024 are available via Zenodo (https://zenodo.org/records/15833452), and data from May 11, 2024 onwards can be obtained from the ISSDC PRADAN portal (https://pradan.issdc.gov.in/al1/). The ULEIS data from the ACE spacecraft are available through NASA’s CDAWeb (https://cdaweb.gsfc.nasa.gov/).